# The distribution of misalignment angles in multipolar planetary nebulae

### v1

@@ 1 Introduction
Most studies attribute each symmetry axis to a pair of jets launched by a binary system progenitor (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024 for a list of a small fraction of all papers; Baan et al. 2021 present an alternative scenario based on fallback material). In most cases, the binary system progenitor experiences a common envelope evolution of a main sequence star orbiting inside the envelope of an asymptotic giant branch (AGB) star; a red giant branch (RGB) star as the primary star is also possible (e.g., Hillwig et al. 2017; Sahai et al. 2017; Jones et al. 2020, 2022, 2023).
@@ 2 The sample of multipolar PNe
In Tables 1 and 2, we list the 40 PNe and 5 pre-PNe, respectively, for which we could find two or more well-defined symmetry axes. In the second column, we give the angle between the axes in each PN (in degrees), referencing an image in the last column.
@@
Figure 3 presents the PN J320, where we identify three symmetry axes, two with pairs of clumps and one with a pair of bubbles. We present this PN to demonstrate a case where the location of the poles is not well defined and where the symmetry axes that we draw by the poles miss the center somewhat. The uncertainties in the projected angles are about $2^{\circ}$. Because we study the cumulative distribution function (Section 3), such uncertainties have negligible influence on our conclusions.
@@ 4 Summary and discussion
From the 40 PNe in our analysis, only M 2-46 is in the catalog of PNe with central binary systems built and maintained by David Jones (Jones & Boffin, 2017; Boffin & Jones, 2019; Jones, 2024). M 2-46 is different from most other multipolar PNe in that one pair is much smaller than the other and inside it (Figure 1).
@@
Consider a protrusion at a distance $h_{c}$ from the equatorial plane. A typical value is about half the scale height $h_{c}\approx 0.5l_{P}\simeq 0.15R_{\rm G}$, whereas we take $l_{P}\simeq 0.3R_{\rm G}$ for the scale height of AGB stars with radius $R_{\rm G}$. The random component due to the convective cells has two terms. The first is due to the orbital motion and the distance of the convective protrusion from the equatorial plane (the orbit) and has a radial (negative or positive) direction
@@
where we take the $z$ axis along the orbital angular momentum direction. The crude estimates of this simple scenario give the stochastic variations, equations (1) and (2), to be several tens of percent of the constant component of the specific angular momentum (equation 3); this might account for our findings of the best angle distribution to fit the observed cumulative distribution function. Future studies should explore the properties of the scenario we propose here and examine other scenarios, like accretion at the CEE’s termination or even the star’s tidal destruction at the core.
@@ Acknowledgements
We thank Gunter Cibis and Martin Guerrero for their valuable comments. We heavily used the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick, 2006) https://faculty.washington.edu/balick/PNIC/. A grant from the Pazy Foundation supported this research.

### v2

@@ 1 Introduction
Most studies attribute each symmetry axis to a pair of jets launched by a binary system progenitor (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Ablimit 2024; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024 for a list of a small fraction of all papers; Baan et al. 2021 present an alternative scenario based on fallback material). In most cases, the binary system progenitor experiences a common envelope evolution of a main sequence star orbiting inside the envelope of an asymptotic giant branch (AGB) star; a red giant branch (RGB) star as the primary star is also possible (e.g., Hillwig et al. 2017; Sahai et al. 2017; Jones et al. 2020, 2022, 2023).
@@ 2 The sample of multipolar PNe
In Tables 1 and 2, we list the 40 PNe and 5 pre-PNe, respectively, for which we could find two or more well-defined symmetry axes. In the second column, we give the angle between the axes in each PN (in degrees), referencing an image in the last column. The sample includes all multipolar PNe and pre-PNe we could find in the literature and from private communications (e.g., the acknowledgment section and searching the Astrophysics Data System-ADS), for which the images were clear enough to define two or more axes.
@@
Figure 3 presents the PN J320, where we identify three symmetry axes, two with pairs of clumps and one with a pair of bubbles. We present this PN to demonstrate a case where the location of the poles is not well defined and where the symmetry axes that we draw by the poles miss the center somewhat. The symmetry lines we draw in our sample can fall in a typical section of about $5^{\circ}$, namely, the typical uncertainties in the projected angles are about ±$\pm 3^{\circ}$; in some cases smaller, in some larger. However, because we study the cumulative distribution function (Section 3), such uncertainties have negligible influence on our conclusions.
@@ 4 Summary and discussion
From the 40 PNe in our analysis, only M 2-46 is in the catalog of PNe with central binary systems built and maintained by David Jones (Jones & Boffin, 2017; Boffin & Jones, 2019; Jones, 2024); see also Chen et al. (2025). M 2-46 is different from most other multipolar PNe in that one pair is much smaller than the other and inside it (Figure 1). Recently, Chen et al. (2025) find variability with a period of 0.0490 in Kn 26, suggesting it is a close binary system. The two pairs of Kn 26 differ substantially in size and shape, similarly to the two pairs of M 2-46.
@@
Consider a protrusion at a distance $h_{c}$ from the equatorial plane. A typical value is about half the typical size of a convective cell. In the mixing length theory, the size of a convective cell is about the pressure scale height, $l_{P}$, or somewhat (up to a factor of about 2) larger. In AGB stars, $l_{P}\simeq 0.3R_{\rm G}$ where $R_{\rm G}$ is the AGB radius. Therefore, we take $h_{c}\approx 0.5l_{P}\simeq 0.15R_{\rm G}$. The random component due to the convective cells has two terms. The first is due to the orbital motion and the distance of the convective protrusion from the equatorial plane (the orbit) and has a radial (negative or positive) direction
@@
where we take the $z$ axis along the orbital angular momentum direction. The crude estimates of this simple scenario give the stochastic variations, equations (1) and (2), to be several tens of percent of the constant component of the specific angular momentum (equation 3); this might account for our findings of the best angle distribution to fit the observed cumulative distribution function.
The stochastic variations that the convection motion introduces change on a time scale of the convective turnover time, $\simeq\tau_{\rm conv}\equiv l_{c}/v_{\rm conv}$. In AGB stars, the convection velocity is a large fraction of the sound speed, which is a large fraction of the Keplerian speed on the surface of the AGB star, $v_{\rm Kep}$. Therefore, the stochastic variation occurs on a time scale similar to the orbital period of a test particle on the surface of the AGB star or shorter, i.e., $\tau_{\rm var}\simeq\tau_{\rm conv}\simeq 0.3R_{\rm G}/v_{\rm conv}\approx R_{\rm G}/v_{\rm Kep}<2\pi R_{\rm G}/v_{\rm Kep}=\tau_{\rm Kep}$. The conclusion is that the convective motion can cause changes in the angular momentum axis in consecutive periastron passages, with a time-lapse of months to years. At the PN phase that occurs hundreds to thousands of years later, the jets will be observed to be coeval, like in the ‘starfish PNe’, e.g., M1-37 (e.g., Sahai 2000). The stochastic convective motion will also cause angular momentum fluctuation in cases where two or more jet-launching episodes are many years (many orbital periods) apart, as in KjPn8.
Hsia et al. (2014) suggest that the lobes result from the interactions between later-developed fast winds and previously ejected asymptotic giant branch winds. However, they do not specify the cause of the change in the directions of the symmetry axes. Steffen et al. (2013) simulate the interaction of a fast wind with a filamentary AGB shell. The filamentary structure is a stochastic one that is not point-symmetric. Namely, there are no pairs of opposite structures. They attribute the filamentary structure to instability or convective motion but not jets. Indeed, the multipolar structures they obtain are not point-symmetric. They do not obtain the structures of most of the PNe we study here, which are point-symmetric.
Future studies should explore the properties of the scenario we propose here and examine other scenarios, such as accretion at the CEE’s termination or the star’s tidal destruction at the core.
@@ Acknowledgements
We thank Gunter Cibis, Martin Guerrero, and an anonymous referee for their valuable comments. We heavily used the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick, 2006) https://faculty.washington.edu/balick/PNIC/. A grant from the Pazy Foundation supported this research.